The National Institute of Health conducted a study to test male birth control pills by using two groups of mice as the study participants. Group A received a drug called Nokids which cost $20,000 and Group B received a drug called Noparents which cost $5,000. The administrative cost to administer each drug is $500 a day. Group A received the drug for 10 days and Group B received the drug for 14 days. What is the incremental cost of Nokids? While conducting the experiment the researchers learned that both medications were contaminated with a poisonous substance. Group A had a survival rate (incremental effectiveness) of 18 days while Group B had a survival rate (incremental effectiveness) of 14 days. What is the incremental cost of Nokids?

(a) $22,500

(b) $18,000

(c) $12,500

(d) $10,000
Answer #1

Cost of Nokids(Group A): $ 20,000

Cost of Noparents(Group B): $5,000

Administrative cost (per day) each drug:$500

No. Of days Group A recieved the drug Nokid =10 days

So, Administrative cost fro Group A drug =10*500=$5,000

No. Of days Group B received the Drug Noparents: 14 days

Total administrative cost of Group B drug:500*14=$7000

Total cost for Group A drug=$20,000+5000=$25,000

Total cost of Group B drug=$5000+7000=$12,000

Here, On calculation of cost ,the effectiveness rate upto which drug shows survival rate is not relevant .As we are calculating cost that we incurr and not the sucess of drug in no. of days.

So incremental cost for Group A drug =$25,000-$12,000(Total cost of Nokids -Total cost of Noparents)

=$12,000

Thus incremental cost is $12,000 ,(no options in the question are appropriate for the answer).

Incremental cost in simple language is extra cost that we incurr in comparison to others for the supply of services or goods.

So,here cost incurred is cost of Drugs and administrative cost to make them delivered and not include survival rate days.